# SoK: Watermarking for AI-Generated Content

### v1

@@ 3.4 Unforgeability
In more detail, unforgeability says that it is computationally infeasible for an attacker without knowledge of an embedding key to produce watermarked content that was not output verbatim by the watermarked model. Note that, while unforgeability is related to the false positive rate, the latter says nothing about the success of a dedicated attacker intentionally trying to get their content flagged.
The first thing to observe about unforgeability is that it is fundamentally incompatible with robustness. If a detector is robust, then it is possible to generate a response from the model, modify it slightly to make it problematic, and then falsely attribute the problematic content to the watermarked model.
However, this impossibility can be circumvented by simply using separate algorithms for detection (which should be robust) and attribution (which should be unforgeable). It is possible to construct a single watermark generation algorithm such that later, at detection/attribution time, one can choose whether to use the detection or attribution algorithm depending on whether one desires robustness or unforgeability. Therefore, we keep the attribution algorithm and key ($\mathsf{Attribute}$ and $\mathsf{ak}$) separate from the detection algorithm and key ($\mathsf{Detect}$ and $\mathsf{dtk}$).
@@ 3.5 Support for Embedding Messages
If the watermark generation algorithm can embed messages in the the content, then we say that it is a multi-bit watermark. Such watermarks are crucial for meeting the increasing demands of customization in various applications of generative models. For instance, embedding detailed metadata, such as model version and user-specific information, enables efficient traceability of the text’s origin across multiple models and users. These capabilities are particularly vital in large-scale deployments.
@@ 5.2.1 Evasion Attacks
    Depending on the threat model, these attacks can make use of automated optimization procedures. For example, an attacker with white-box access to a watermark detector could optimize adversarial noise that fools the detector [67]. An attacker could also first “steal” the watermarking key [68] and then make targeted edits to remove the watermark.
@@ 6 Recent Representative Watermark Schemes
In this section, we present several representative works in the field to illustrate how watermarking schemes operate. We also discuss the strengths and limitations of different methods to provide a balanced understanding of their effectiveness.
@@ 6.1 Text Watermarks
Numerous studies have explored variations and improvements to Green-Red Watermarking, including [102, 103, 104, 105, 106], among others.
@@ 6.3 Video and Audio Watermarks
Compared to the advancements in text and image watermarking, relatively few studies focus on video and audio watermarks specifically designed for generative AI. For video watermarking, existing methods often rely on traditional video watermarking techniques [127, 128] or apply post-processing watermarks to individual frames [120, 129], similar to image watermarking. In the case of audio watermarking, many approaches embed watermarks into synthetic audio using a trained/fine-tuned watermark encoder, primarily targeting text-to-speech generative models [130, 31, 131, 132, 32, 133, 134].
@@ 7 Open Problems and Discussion
Robustness and Unforgeable Public Attribution. Who should have the right to detect watermarked text? In the most democratized case, the public attribution setting, watermarks can function analogously to digital signatures where a private entity is allowed to embed watermarks and anyone can detect their presence. From a technical perspective, it is possible to construct watermarks with unforgeable public attribution: the open accessibility of the detector does not introduce forgery attacks. However, present constructions either lack strong robustness or concrete efficiency. It remains open whether schemes with public attribution and strong robustness can be efficiently instantiated or if more powerful robustness notions can coexist with public attribution in theory.
@@
Other Applications of Watermarking.. Watermarking extends beyond content traceability and offers applications in various aspects of AI model and data governance. For instance:
@@
Policy Challenges and the Future of Watermarking. For watermarking to be universally effective in identifying AI-generated content, it requires widespread adoption by all organizations offering generative AI services. A unified watermarking scheme would enable comprehensive detection of AI-generated content across platforms. Without such standardization, detection efforts are likely to be fragmented, with model owners only capable of identifying outputs generated by their own systems. Achieving this level of coordination poses significant challenges. Open collaboration among companies, coupled with governmental legislation and global policy frameworks, is necessary to establish and enforce universal watermarking standards. Such measures would ensure that AI-generated content remains traceable, thereby addressing concerns related to misinformation, intellectual property infringement, and ethical accountability. The future of watermarking depends not only on technological advancements but also on cohesive policy efforts across international boundaries. It is important that these policy efforts are aligned with state-of-the-art watermarking schemes, cognizant of their differences, capabilities, and limitations [53].

### v3

@@ 1 Introduction
1.1 Methodology for Systematizing Knowledge
To construct this SoK, we conducted an extensive review of over 100 peer-reviewed papers published from the early 2000s to the 2020s. Our sources include top venues in security (e.g., USENIX Security, IEEE S&P), AI (e.g., NeurIPS, ICML), and multimedia (e.g., CVPR, ACM Multimedia), as well as relevant arXiv preprints reflecting the fast-paced development of the field. We further incorporated key regulatory and industry documents, such as the EU AI Act and C2PA specifications. Our selection criteria prioritized works by their relevance to GenAI watermarking across various modalities, contribution to definitions, threat modeling, evaluation, or technical advances, and identification of unresolved issues and future directions. This structured approach enabled impartial and accurate representation of diverse perspectives and established principles within the surveyed literature.
1.2 Comparison with Prior Surveys
While watermarking has a long history, its application to GenAI content is a relatively recent and rapidly developing field. Several surveys have touched upon related aspects. Traditional digital watermarking surveys (e.g., [35]) provide foundational concepts but predate modern GenAI. Surveys on deep learning watermarking often focus on model watermarking for ownership protection (e.g., [36, 37]), which has different goals and technical challenges than content watermarking. Amrit and Singh [38] provide a broad overview of watermarking in AI but with limited focus on generative models specifically.
In contrast, our SoK offers a timely and in-depth analysis focused specifically on watermarking outputs of GenAI models. Our key contributions are: (1) Scope: Coverage across multiple modalities—text, image, audio, and video. (2) Depth: Formal definitions, structured threat models, and evaluation criteria. (3) Timeliness: Inclusion of recent advances, such as undetectable and provable watermarking schemes. Together, these elements position our paper as a comprehensive and authoritative resource on watermarking in the GenAI era.
@@ 3.4 Unforgeability
In more detail, one natural way to define unforgeability is to require that it be computationally infeasible for an attacker without knowledge of an embedding key to produce watermarked content that was not output verbatim by the watermarked model. Note that, while unforgeability is related to the false positive rate, the latter says nothing about the success of a dedicated attacker intentionally trying to get their content flagged.
The first thing to observe is that this definition of unforgeability is fundamentally incompatible with robustness. If a detector is robust, then it is possible to generate a response from the model, modify it slightly to make it problematic, and then falsely attribute the problematic content to the watermarked model, which could be undesirable in some cases.
However, this incompatibility can be circumvented by simply using separate algorithms for detection (which should be robust) and attribution (which should be unforgeable). It is possible to construct a single watermark generation algorithm such that later, at detection/attribution time, one can choose whether to use the detection or attribution algorithm depending on whether one desires robustness or unforgeability. Therefore, we keep the attribution algorithm and key ($\mathsf{Attribute}$ and $\mathsf{ak}$) separate from the detection algorithm and key ($\mathsf{Detect}$ and $\mathsf{dtk}$). An alternative way would be to modify the unforgeability requirement so that it is computationally infeasible for an attacker without knowledge of an embedding key to produce watermarked content that is dissimilar — rather than verbatim — from the content generated by the watermarked model. Here, we focus on the first approach and refer the reader to [75] for more discussion on the second.
@@ 3.5 Support for Embedding Messages
If the watermark generation algorithm can embed messages in the content, then we say that it is a multi-bit watermark. Such watermarks are crucial for meeting the increasing demands of customization in various applications of generative models. For instance, embedding detailed metadata, such as model version and user-specific information, enables efficient traceability of the text’s origin across multiple models and users. These capabilities are particularly vital in large-scale deployments.
@@ 5.2.1 Evasion Attacks
    Depending on the threat model, these attacks can make use of automated optimization procedures. For example, an attacker with white-box access to a watermark detector could optimize adversarial noise that fools the detector [70]. An attacker could also first “steal” or “estimate” the watermarking pattern [71, 77] and then make targeted edits to remove the watermark.
@@ 6 Recent Representative Watermark Schemes
In this section, we present several representative works in the field to illustrate how watermarking schemes operate. We discuss the strengths and limitations of different methods to provide a balanced understanding of their effectiveness. We also connect these representative works with the desired properties of a watermarking scheme discussed in Section 3. An approximate conceptual comparison is shown in Table 1.
@@ 6.1 Text Watermarks
Numerous studies have explored variations and improvements to Green-Red Watermarking, including [106, 107, 108, 109, 110, 111], among others.
Method Quality FPR Robustness Unforgeability Messages Efficiency
Text Watermarking
Green-Red WM [26, 27]
Gumbel WM [25]
Undetectable WM [28]
Pseudorandom Codes [73]
Semantic Sentence WM [112]
Image Watermarking
Stable Signature [113]
Tree-Ring WM [29]
Gaussian Shading WM [69]
PRC Watermark [30]
Table 1: Qualitative comparison of representative watermarking methods across key evaluation criteria. Darker circles indicate stronger performance, offering an intuitive view of trade-offs without numerical scores.
@@ 6.3 Video and Audio Watermarks
Compared to the advancements in text and image watermarking, relatively few studies focus on video and audio watermarks specifically designed for generative AI. For video watermarking, existing methods often rely on traditional video watermarking techniques [132, 133] or apply post-processing watermarks to individual frames [126, 134], similar to image watermarking. In the case of audio watermarking, many approaches embed watermarks into synthetic audio using a trained/fine-tuned watermark encoder, primarily targeting text-to-speech generative models [135, 31, 136, 137, 32, 138, 139]. Looking forward, watermarking for generative media could benefit from designs tailored to any-to-any generation pipelines [140], including cross-modal generation [141] and even other modalities such as tabular data generation [142].
@@ 7 Open Problems and Discussion
Robustness and Unforgeable Public Attribution. Who should have the right to detect watermarked content? In the most democratized case, the public attribution setting, watermarks can function analogously to digital signatures where a private entity is allowed to embed watermarks and anyone can detect their presence. From a technical perspective, it is possible to construct watermarks with unforgeable public attribution: the open accessibility of the detector does not introduce forgery attacks. However, present constructions either lack strong robustness or concrete efficiency. It remains open whether schemes with public attribution and strong robustness can be efficiently instantiated or if more powerful robustness notions can coexist with public attribution in theory.
@@
Other Applications of Watermarking. Watermarking extends beyond content traceability and offers applications in various aspects of AI model and data governance. For instance:
@@
Policy Challenges and the Future of Watermarking. For watermarking to be universally effective in identifying AI-generated content, it requires widespread adoption by all organizations offering generative AI services. A unified watermarking scheme would enable comprehensive detection of AI-generated content across platforms. Without such standardization, detection efforts are likely to be fragmented, with model owners only capable of identifying outputs generated by their own systems. Achieving this level of coordination poses significant challenges. A number of measures may be necessary to establish and enforce universal watermarking standards – including open collaboration among companies coupled with governmental legislation and global policy frameworks. Such measures would ensure that AI-generated content remains traceable, thereby addressing concerns related to misinformation, intellectual property infringement, and ethical accountability. The future of watermarking depends not only on technological advancements but also on cohesive policy efforts across international boundaries. It is important that these policy efforts are aligned with state-of-the-art watermarking schemes, cognizant of their differences, capabilities, and limitations [56].